# FissionVAE: Federated Non-IID Image Generation with Latent Space and Decoder Decomposition The source code of FissionVAE is publicly available at: https://github.com/Rand2AI/FissionVAE

### v1

@@ 4.2 Results and Analysis
We conduct four sets of experiments on the Mixed MNIST and CHARM dataset. In the first experiment, we evaluate the overall generative performance of the proposed VAE architectures in both federated and centralized settings. The second experiment focuses on the hierarchical VAE’s ability to control the generation process through explicit definition of the prior distributions for $z_{1}$. The hierarchical model allows for multiple generation pathways. For instance, one pathway involves sampling $z_{2}$, passing it through the decoder $p(z_{1}|z_{2})$, and then generating the final image with $p(x|z_{1})$. Alternatively, we could directly sample $z_{1}$ and generate the image through $p(x|z_{1})$. We compare the performance of these different generation pathways to assess their impact on the quality of generated images. Following the discussion on how $z_{1}$ affects the generation performance, we explore various strategies for defining $z_{1}$’s prior distribution and assess the quality of generated images following the pathway $z_{2}\rightarrow z_{1}\rightarrow x$. Lastly, we showcase the use of heterogeneous decoder architectures in the proposed federated VAE and the effects of pixel modeling methods.
@@ 4.2.1 Overall Performance
The overall performance of the proposed FissionVAE models is summarized in Table 3, and generated examples are shown in Fig. 8. As a baseline, a Deep Convolutional GAN (DCGAN) [21] trained via FedGAN [22] is used for comparison. Since GAN does not directly model the likelihood of data, NLL is not evaluated for FedGAN. Also, FedGAN on CHARM suffers from sever mode collapse (shown in Fig. 9), therefore performance evaluation is not available on this dataset. Notably, the performance of all models on the CHARM dataset was less robust compared to the Mixed MNIST dataset. This discrepancy arises because the CHARM dataset, encompassing colorful images from diverse domains, presents a more complex and realistic federated learning scenario. The dataset’s diversity, coupled with a lower local data availability and participation rate among clients, poses greater challenges to federated generative models.
Figure 9: Example of Collapesed Models. First Row: DCGAN suffers from severe mode collapse on CHARM in both centralized and federated settings. Second Row: When the “symmetrical” prior encoding is used on the CHARM dataset, priors that are further away from the origin leads to complete collapse regard. Remote sensing group is assigned a prior mean of 1 on the left, on the right 3 is assigned to this group, resulting in complete collapse.
Latent Space Decoupling vs Decoder Branches Across the datasets and learning paradigm, we observe the trend that adopting latent space decoupling or group-specific decoder branches improves the quality of generated images with lower FID and higher IS. We can also tell from Table 3 that the creating decoder branches alone brings more performance improvement compared to sole latent space decoupling, showing that mixing decoders trained on non-IID data can severely affect image generation.
@@
Among the proposed VAE configurations, the model FissionVAE+L (L for decoupled latent space) demonstrates a slight improvement over the baseline FissionVAE. This model adapts the latent space to the non-IID nature of nature of federated data by partitioning it according to client groups. This partitioning not only informs the decoder of the expected data characteristics more effectively but also helps to prevent the overlap of data representations from different domains. The effect of latent space decoupling can be also observed in Fig. 8. Compared to the baseline FedVAE, FissionVAE+L allows for directly sampling the latent codes specific to each client group, showing more intra-group difference in generated images. However, as the decoders in FissionVAE+L are mixed during the aggregation phase, this mixing results in noticeable artifacts that resemble overlapping images from different domains, such as creating contours of faces for latent codes sampled from marine vessels’ distribution.
@@
FissionVAE+D (D for decoder branches) greatly alleviates the problem of overlapping features. The FissionVAE+D architecture consists of a unified encoder and decoder branches for individual client groups. The unified encoder acts as a routing mechanism similar to that in a Mixture of Expert (MoE) framework, directing data to the appropriate latent distributions based on its domain. During model aggregation, these decoders among different client groups remain distinct, ensuring that textural features from different domains do not mix, thus producing cleaner and more discernible images as shown in Fig 8.
@@
FissionVAE+L+D is the architecture for combining both latent space decoupling and decoder branches, Table 3 shows that there is little gain on the Mixed MNIST dataset while the generative performance is further improved on the CHARM dataset. Enforcing latent space decoupling on FissionVAE+D even lowers the performance on FID. We argue that the main reason for this is the number of client groups. As there are only two groups in Mixed MNIST, decoder branches alone is sufficient for the VAE to differentiate samples from different groups. However, as the number of client groups increases, explicit latent space decoupling provides more direct signal to the VAE about the intra-group difference, allowing the model to better capture the data distribution for different groups. In Fig. 8 it can be observed that images generated by FissionVAE+L+D are sharper than the ones generated by FissionVAE+D.
@@
Hierarchical FissionVAE As discussed in Section III, here we consider a hierarchical VAE with two levels of latent variable. In Table 3, the architecture FissionVAE+H+L+D performs the best on the CHARM dataset and falls behind its non-hierarchical counterpart on the Mixed MNIST dataset. The hierarchical VAE employs multiple levels of latent representations, which refines the model’s ability to capture and reconstruct complex data distributions more faithfully. The performance degradation on simpler datasets like Mixed MNIST suggests that the hierarchical approach might introduce unnecessary redundancy without proportional gains in performance.
Table 4: Definition of Means for the Prior Distribution of $z_{1}$
Approach Definition of Means Example of Two Client Groups
identical $\bm{\mu}^{(j)}=\bm{0}$ [0, 0, 0, 0], [0, 0, 0, 0]
one-hot $\mu_{j}^{(j)}=1$ [0, 1, 0, 0], [0, 0, 1, 0]
symmetrical $\mu_{i}^{(j)}=\lceil\frac{j}{2}\rceil$ (j is even) or $-\lceil\frac{j}{2}\rceil$ (j is odd) [1, 1, 1, 1], [-1, -1, -1, -1]
random $\mu_{i}^{(j)}\sim N(0,1)$ [0.9, 0.3, 1.6, -0.5], [1.5, -1.8, 0.1, 0.6]
wave $\mu_{i}^{(j)}=1,i\in\left[j\lfloor\frac{dim(z_{1})}{k}\rfloor,(j+1)\lfloor\frac{dim(z_{1})}{k}\rfloor\right)$ [1, 1, 0, 0], [0, 0, 1, 1]
learnable learn from $D_{\text{KL}}(q_{\phi}(z_{1}|x)||p_{\theta}(z_{1}|z_{2}))$
Elements that are not defined above are 0’s by default. $\mu_{i}^{(j)}$ denotes the $i$-th dimension of the mean for $z_{1}$ from the client group $j$.
$k$ is the total number of client groups.
Table 5: Evaluation of Generation Performance with $z_{1}$ Priors
@@
4.2.2 Decoupling the Prior of $z_{1}$
As shown in Table 3, explicitly decoupling the latent space for different client groups can improve the VAE’s ability to generate images that better reflect the true data distribution. We explore various approaches to define the prior for the latent distribution, all of which are assumed to be multivariate Gaussians with customizable means and the identity matrix as their fixed covariance matrices. The definition of investigated priors are listed in Table 4, and their evaluation results are presented in Table 5.
For the non-hierarchical VAE, there is a single latent variable $z_{1}$. In the two-level hierarchical VAE, we control only the first-level latent variable $z_{1}$, while the prior distribution for $z_{2}$ remains a standard normal distribution $N(0,1)$. Regardless of the hierarchical architecture, the ”identical” approach in Tables 4 and 5 serves as a baseline with no latent space decoupling, where the prior distributions are identical for all client groups and are defined as the normal distribution. The ”one-hot” approach assigns one-hot encoded vectors as the means of $z_{1}$ across different client groups. In the ”symmetrical” approach, the means of $z_{1}$ are set to positive and negative integers symmetrical to zero. For instance, the means for groups 1 and 2 are ±$\pm 1$, respectively, and for groups 3 and 4, they are ±$\pm 2$, and so on. The ”random” approach involves sampling random vectors from a normal distribution as the means of $z_{1}$ for each client group. The ”wave” strategy extends the one-hot encoding by including more ones. Specifically, the dimension of $z_{1}$ is evenly divided based on the number of client groups. The $j$-th division is set to 1 for the $j$-th client group, with the rest of the dimensions being zero. Finally, the ”learnable” approach is unique to hierarchical VAEs, arising from the prior matching term between the encoded and decoded $z_{1}$ in Equation 6. This allows the model to learn to align the prior for each client group, but we lose the ability to directly sample from $z_{1}$ for the controlled generation process. In contrast, other approaches also need to optimize the KL divergence between the encoder’s output $q_{\phi}(z1|x)$ and the defined prior $p(z_{1})$ in addition to the prior matching divergence.
From Table 5, we observe that when predefined priors are enforced on $z_{1}$, hierarchical FissionVAE generally underperforms compared to the non-hierarchical FissionVAE. This performance disparity may stem from the increased uncertainty in the latent space introduced by the hierarchical structure. Specifically, the presence of multiple latent layers in hierarchical VAEs introduces more stochasticity, which, coupled with limited local training samples and non-IID global data, causes greater variance in latent representations across different clients. This variance often results in degraded generation quality. Notably, the hierarchical VAE achieves optimal performance with a learnable approach for $z_{1}$, where no fixed prior distribution is enforced. This flexibility allows the model to adapt the latent space dynamically, enhancing generation quality. However, implementing a similar learnable approach in non-hierarchical VAEs is impractical as the ELBO formulation in Equation 2 requires the latent distribution to match the predefined prior.
On the Mixed MNIST dataset, employing the identical approach, where a standard normal distribution is used across all client groups, yields the lowest FID for the non-hierarchical VAE but performs poorly on the CHARM dataset. This outcome suggests that for simpler datasets with only two client groups, creating decoder branches alone is adequate for generating distinct samples from different data distributions. However, as the number of client groups increases, a more explicit latent encoding can better guide the model to capture complex data distributions. The identical approach shows diminished performance on both datasets in the hierarchical VAE context. This issue likely arises from the inference hierarchy of the latent variables. Since $z_{2}$ is estimated by the encoder $q_{\phi}(z_{2}|z_{1})$, sampling $z_{1}$ from the same distribution across different client groups can introduce contradictions in the latent space.
@@ 4.2.2 Decoupling the Prior of $z_{1}$
Table 7: Evaluation of Image Generation from Different Pathways in FissionVAE+H+L+D
Generation Pathway Mixed MNIST CHARM
FID $\downarrow$ IS $\uparrow$ FID $\downarrow$ IS $\uparrow$
$q(z_{1})\rightarrow q_{\phi}(z_{2}|z_{1})\rightarrow p_{\theta}(z_{1}|z_{2})\rightarrow p_{\theta}(x|z_{1})$ 93.79 2.43 141.97 2.02
$q(z_{1})\rightarrow p_{\theta}(x|z_{1})$ 124.76 2.41 144.26 2.05
$p(z_{2})\rightarrow p_{\theta}(z_{1}|z_{2})\rightarrow p_{\theta}(x|z_{1})$ 53.68 2.94 118.56 2.24
The symmetrical approach, where the mean of $p(z_{1})$ is set as the positive or negative integer of the group index, can lead to severe model divergence on the CHARM dataset, shown in Figure 9. As the number of client groups increases, this strategy results in means for $p(z_{1})$ that far exceed the initialized values of the neural network, challenging the model’s ability to converge, particularly for client groups with large means. The one-hot and random approaches show comparative performance but it falls behind the wave approach in terms of the consistency of lower FID on both datasets. Both the one-hot and random approaches exhibit comparative performance, yet they fall behind the wave approach in terms of consistently lower FID across both datasets. The wave approach offers a clearer distinction between the prior distributions for different client groups. Moreover, since all non-zero elements in the wave approach are 1’s, it avoids the out-of-range value problem observed with the symmetrical approach.
4.2.3 Generation Pathways in Hierarchical FissionVAE
Following the discussion on decoupling the latent space of $z_{1}$ by defining prior distributions for individual client groups, we can now sample from either $q(z_{1})$ or $p(z_{2})$ to produce generated samples. There are two possible generation pathways for $z_{1}$. The first involves treating random samples from the prior distribution as outputs from the encoder $q_{\phi}(z_{1}|x)$, which are then passed to the encoder $q_{\phi}(z_{2}|z_{1})$ for further processing before the hierarchical decoder outputs the final generated samples. The second pathway treats $z_{1}$ samples as outputs from the decoder $p_{\theta}(z_{1}|z_{2})$, directly generating image samples from the final decoder $p_{\theta}(x|z_{1})$. The pathway for $z_{2}$ follows that of non-hierarchical VAEs, involving only the decoders in the generation process. We list the discussed pathways and their evaluation results in Table 7, and the prior distributions of $z_{1}$ are defined with the wave approach in Table 4.
Figure 10: When the prior distributions of the MNIST and FashionMNIST groups are swapped, the generation pathway $q(z_{1})\rightarrow q_{\phi}(z_{2}|z_{1})\rightarrow p_{\theta}(z_{1}|z_{2})\rightarrow p_{\theta}(x|z_{1})$ leads to sever mode collapse, suggesting potential group-level privacy preserving through protected prior distribution.
@@ 4.2.3 Generation Pathways in Hierarchical FissionVAE
As indicated in Table 7, the conventional generation pathway from $z_{2}$ consistently outperforms those from $z_{1}$. Including the encoder in the pathway from $z_{1}$ allows for mapping random inputs to more meaningful representations for subsequent decoders, enhancing performance compared to the direct generation pathway using $p_{\theta}(x|z_{1})$. Notably, on the Mixed MNIST dataset, swapping the prior distributions of the two client groups in the encoder-inclusive pathway leads to significant mode collapse, shown in Figure 10. This suggests that the group-level privacy may be preserved by maintaining the confidentiality of prior distributions. This strategy ensures that high-quality samples are generated only when the correct prior distribution is used, while mismatched distributions yield unrecognizable outputs. This phenomenon is more pronounced in both hierarchical and non-hierarchical FissionVAEs on the Mixed MNIST dataset than on the CHARM dataset, likely due to the simpler, more uniform nature of the Mixed MNIST data compared to the diverse and colorful image types in CHARM, which pose greater challenges in satisfying complex latent distribution constraints.
@@ 4.2.4 Heterogeneous Decoders in FissionVAE
As discussed in Section III, the distinct separation and aggregation of decoders tailored to specific client groups allow for the use of heterogeneous architectures in FissionVAE. The Mixed MNIST dataset, with its relatively simple and grayscale colors, can be generated from both fully connected (MLP) and convolutional layers. In contrast, the more complex and colorful images in the CHARM dataset predominantly require convolutional layers for effective generation.
Table 6 details the performance evaluation of various decoder architectures. The term ‘homogeneous’ refers to identical architectural configurations across all decoder branches, namely a three-layer MLP for each decoder modules. In the ‘Deeper MLP’ configuration, we add two additional fully connected layers to both $p_{\theta}(z_{1}|z_{2})$ and $p_{\theta}(x|z_{1})$. Meanwhile, we completely replace the decoder $p_{\theta}(x|z_{1})$ from MLP to a series of transpose convolution layers in the ‘Deeper MLP + Conv’ configuration. The results indicate a gradual reduction in overall FID scores as the decoder architecture becomes more heterogeneous. However, the integration of convolutional layers does not improve generation performance over the MLP models, underscoring that while heterogeneous architectures are feasible, they can disrupt the convergence of the VAE due to mismatches in architecture and the model’s weight space.
Table 8: Comparison between Discretized Mixture of Logistics and Gaussians as the reconstruction loss.
Model Reconstruction Loss FID $\downarrow$ IS $\uparrow$
FissionVAE+L+D (wave) DLML 85.12 3.54
Gaussian 109.10 2.27
FissionVAE+H+L+D (wave) DLML 82.35 3.91
Gaussian 118.56 2.24
Figure 11: Comparison Between the Discretized Logistic Mixture Likelihood and the Gaussians as the Reconstruction Loss. First row reconstructs test images while the second row generates images from random input.
4.3 Modeling the Pixels
The reconstruction term in the ELBO quantifies the likelihood of the original input given the estimated distribution from the decoder. L1 or L2 losses, which measure absolute or squared differences, respectively, do not align well with the probabilistic outputs of VAEs. These metrics fail to capture the underlying data distribution effectively, making them less suitable for VAEs that model the complex variability of image pixels.
For grayscale images in the Mixed MNIST dataset, we follow the convention by modeling the reconstruction term using binary cross-entropy between the decoder output and the original input, which captures the presence or absence of pixel values in binary images. For colorful images, modern VAEs often employ the Discretized Logistic Mixture Likelihood (DLML), which models the pixel values more flexibly by considering a mixture of logistic distributions, thus capturing a wider range of color dynamics. In our implementation of FissionVAE, we consider a simpler Gaussian model for each pixel, assuming a shared global variance across all images.
When comparing these two loss function options, we find that while DLML can produce sharper and more visually appealing images, as reflected in higher inception scores, it also introduces stylized artifacts reminiscent of watercolor artworks, characterized by large, irregular color patches. Consequently, we continue to employ the Gaussian formulation in all our implementations due to its simplicity and more consistent performance across varied image types. The qualitative comparison is shown in Figure 11 and the evaluation results are presented in Table 8.

### v2

@@ 3.2 Results and Analysis
Here we present the following experiments: we first evaluate the overall generative performance of the proposed VAE architectures in both federated and centralized settings, then we explore strategies for encoding the prior distribution $p(z_{1})$, and lastly we showcase the use of heterogeneous decoder architectures in our FissionVAEs. For experiments investigating different generation pathways of hierarchical VAEs and the effect of reconstruction losses, please refer to our supplementary material.
@@ 3.2.1 Overall Performance
The overall performance of the proposed FissionVAE models is summarized in Table 1, and generated examples are shown in Fig. 11. In addition to the FedVAE baseline, a Deep Convolutional GAN (DCGAN) Radford et al. (2016) trained via FedGAN Rasouli et al. (2020) is used for comparison. Since GAN does not directly model the likelihood of data, NLL is not evaluated for FedGAN. Also, FedGAN on CHARM suffers from severe mode collapse, therefore performance evaluation is not available on this dataset. Notably, the performance of all models on the CHARM dataset is less robust compared to the Mixed MNIST dataset. This discrepancy arises because the CHARM dataset, encompassing RGB images from diverse domains, presents a more complex and realistic federated learning scenario. The dataset’s diversity, coupled with a lower local data availability and participation rate among clients, poses greater challenges to federated generative models.
Latent Space Decoupling vs Decoder Branches As shown in Table 1, both latent space decoupling and group-specific decoder branches improve image quality, resulting in lower FID and higher IS. However, decoder branches alone yield greater performance improvements than latent space decoupling, highlighting the negative impact of mixing decoders trained on non-IID data.
@@
Among the configurations, FissionVAE+L shows moderate improvement over the baseline FedVAE by partitioning the latent space according to client groups. This partitioning not only informs the decoder of the expected data characteristics more effectively but also helps to prevent the overlap of data representations from different domains. The effect of latent space decoupling can be also observed in the qualitative results. In Fig. 11, although FissionVAE+L allows sampling latent codes specific to each group, mixing decoders during aggregation causes artifacts, such as blending features from different domains.
@@
FissionVAE+D greatly alleviates the problem of blended visual features. The FissionVAE+D architecture consists of a unified encoder and decoder branches for individual client groups. The unified encoder acts as a routing mechanism similar to that in a Mixture of Expert (MoE) framework, directing data to the appropriate latent distributions based on its domain. During model aggregation, these decoders among different client groups remain distinct, ensuring that textural features from different domains do not mix, thus producing cleaner and more discernible images as shown in Fig 11.
@@
FissionVAE+L+D is the architecture that combines both latent space decoupling and decoder branches, Table 1 shows that there is less gain on the Mixed MNIST dataset while the generative performance is further improved on the CHARM dataset. Enforcing latent space decoupling on FissionVAE+D even lowers the performance on FID. We argue that the main reason for this is the number of client groups. As there are only two groups in Mixed MNIST, decoder branches alone is sufficient for the VAE to differentiate samples from different groups. However, as the number of client groups increases, explicit latent space decoupling provides more direct signal to the VAE about the intra-group difference, allowing the model to better capture the data distribution for different groups. In Fig. 11 it can be observed that images generated by FissionVAE+L+D are sharper than the ones generated by FissionVAE+D.
@@
Hierarchical FissionVAE As discussed in Section 2, here we consider a hierarchical VAE with two levels of latent variable. In Table 1, the architecture FissionVAE+H+L+D performs the best on the CHARM dataset and falls behind its non-hierarchical counterpart on the Mixed MNIST dataset. The hierarchical VAE employs multiple levels of latent representations, which refines the model’s ability to capture and reconstruct complex data distributions more faithfully. The performance degradation on simpler datasets like Mixed MNIST suggests that the hierarchical approach might introduce unnecessary redundancy without proportional gains in performance.
@@
Table 2: Evaluation of Generation Performance with $z_{1}$ Priors
@@
3.2.2 Decoupling the Prior of $z_{1}$
Explicitly decoupling the latent space for different client groups improves the VAE’s ability to generate images that align with the true data distribution (Table 1). We explore several priors for the latent distribution, modeled as multivariate Gaussians with customizable means and identity covariance matrices and evaluate them in Table 2. Details regarding the formal definition of priors can be found in the supplementary material.
In non-hierarchical VAEs, $z_{1}$ represents the sole latent variable, while in hierarchical VAEs, $z_{1}$ is controlled, with $z_{2}$ following a standard normal distribution $N(0,1)$. Baseline priors are identical across client groups. Other prior variations include one-hot encoding, symmetrical positive and negative integers, random vectors, wave encodings (with grouped 1s in dimensions corresponding to client groups), and a learnable approach unique to hierarchical VAEs. The learnable approach dynamically aligns priors but sacrifices direct sampling from $p(z_{1})$.
Hierarchical FissionVAE often underperforms non-hierarchical variants when predefined priors are used due to increased uncertainty from additional latent layers. However, the learnable approach excels in capturing complex distributions dynamically. In simpler datasets like Mixed MNIST, identical priors suffice, but explicit latent encoding becomes crucial as client group diversity increases, as seen with CHARM.
Among prior definitions, symmetrical priors often lead to divergence on CHARM, as their means may exceed neural network initialization ranges. One-hot and random approaches show comparable results but are less consistent than wave encoding, which clearly distinguishes group priors without out-of-range values.
3.2.3 Group-level Privacy
@@
In the presence of hierarchical VAEs, it is possible to incorporate the encoder $q_{\phi}(z_{2}|z_{1})$ into the generation process, that is, we can first sample the latent code $z_{1}$ from its prior distribution, then feed it to the subsequent encoder $q_{\phi}(z_{2}|z_{1})$ and the decoders $p_{\theta}(z_{1}|z_{2})$ and $p_{\theta}(x|z_{1})$ to obtain the synthesize a generated sample. On the Mixed MNIST dataset, we observe that swapping the prior distributions of the two client groups in the such a generation pathway leads to evident mode collapse, shown in Figure 7. This suggests that the group-level privacy may be preserved by maintaining the confidentiality of prior distributions. This strategy ensures that high-quality samples are generated only when the correct prior distribution is used, while mismatched distributions yield unrecognizable outputs. This phenomenon is more pronounced in both hierarchical and non-hierarchical FissionVAEs on the Mixed MNIST dataset than on the CHARM dataset, likely due to the simpler, more uniform nature of the Mixed MNIST data compared to the diverse and colorful image types in CHARM, which pose greater challenges in satisfying complex latent distribution constraints. Evaluation on other generation pathways are presented in the supplementary material.
Figure 7: In hierarchical FissionVAE, when the prior distribution $p(z_{1})$ of the MNIST and FashionMNIST groups are swapped, the generation pathway $q(z_{1})\rightarrow q_{\phi}(z_{2}|z_{1})\rightarrow p_{\theta}(z_{1}|z_{2})\rightarrow p_{\theta}(x|z_{1})$ leads to sever mode collapse, suggesting potential group-level privacy preserving through protected prior distribution.
@@ 3.2.4 Heterogeneous Decoders in FissionVAE
As discussed in Section 2, the decoupling of decoders for client groups allow for the use of heterogeneous architectures in FissionVAE. The Mixed MNIST dataset, with its relatively simple and grayscale colors, can be generated from both fully connected (MLP) and convolutional layers. In contrast, the more complex and colorful images in the CHARM dataset predominantly require convolutional layers for effective generation.
Table 3 details the performance evaluation of various decoder architectures. The term ’homogeneous’ refers to identical architectural configurations across all decoder branches, namely a three-layer MLP for each decoder modules. In the ‘Deeper MLP’ configuration, we add two additional fully connected layers to both $p_{\theta}(z_{1}|z_{2})$ and $p_{\theta}(x|z_{1})$. Meanwhile, we completely replace the decoder $p_{\theta}(x|z_{1})$ from MLP to a series of transpose convolution layers in the ‘Deeper MLP + Conv’ configuration. The results indicate a gradual reduction in overall FID scores as the decoder architecture becomes more heterogeneous. However, the integration of convolutional layers does not improve generation performance over the MLP models, underscoring that while heterogeneous architectures are feasible, they can disrupt the convergence of the VAE due to mismatches in architecture and the model’s weight space.